OpenStudy (tiffanymak1996):

apply the ratio test to the following series and state what you can conclude from it.

OpenStudy (tiffanymak1996):

\[\sum_{k=1}^{\infty}\frac{ k ^{50} }{ 2^{k} }\]

OpenStudy (tiffanymak1996):

and the ratio test:

OpenStudy (tiffanymak1996):

\[\rho = \lim_{k \rightarrow + \infty}\frac{ u _{k+1} }{u _{k} }\]

OpenStudy (anonymous):

Can you have a go at applying the ratio test yourself?

OpenStudy (tiffanymak1996):

i did, just can't do the limit part

OpenStudy (tiffanymak1996):

the series is suppose to be convergent, when i work till:\[\lim_{k \rightarrow + \infty } \frac{ (k+1)^{50} }{ 2k ^{50} }\] i don't know how to go on.

OpenStudy (tiffanymak1996):

ah, wait, can i use l'hospital rule?

OpenStudy (anonymous):

Ok that's what I got also. I'm going to neglect writing the limit here but remember it's still there, I'm just lazy. If you expanded the numerator it would look something like this: \[\frac{ k^{50} + ... + 1 }{ k ^{50} }\] Where all the ... terms are just multiples of k^n for n between 49 and 1. Now if we divide by k^50 on top and bottom we get: \[\frac{ 1+...+\frac{ 1 }{ k ^{50} } }{ 2 }\] And for k->infinity, all of the ... and 1/k^50 will cancel.

OpenStudy (anonymous):

(I didn't mean cancel, I mean tend to 0)

OpenStudy (anonymous):

So can you see what we will be left with?

OpenStudy (tiffanymak1996):

1/2

OpenStudy (anonymous):

Indeed :)

OpenStudy (tiffanymak1996):

thx!